OpenStudy (anonymous):

Jada is trying to determine the number of t-shirts in inventory before she started her shift. During her shift, she sold 9 t-shirts, which left 19 t-shirts in inventory. -Define a variable. -Write an equation Jada can use to determine the number of t-shirts that were in inventory before she started her shift. - Solve the equation showing your work

OpenStudy (anonymous):

you can get medal if you try

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

s= number of shirts sold before Jada's shift s+9=19 -9 -9 s=10

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

well here you r thxs

OpenStudy (anonymous):

10 is the answer, how ever you chose to solve it.

OpenStudy (anonymous):

yep